OpenStudy (anonymous):

24 is what percent of 32? write the proportion too please!

OpenStudy (anonymous):

(24/32)*100=

OpenStudy (anonymous):

75%

OpenStudy (anonymous):

75% You do 24/32 =x/100 and solve for x to find a percentage of something. 24/32=.75 then multiply .75 by 100 and you get the answer...75%